An X molal solution of a compound in benezene has mole fraction of solute = .2. The value of X is 
 

Grade:9

1 Answers

Himanshu Rathour
5 years ago
Let mole of solute =x
Mole of solvent=y
Wt of solvent=w
Molality=mole of solute/wt of solvent
Mole fraction=mole of solute/mole of solute+mole of solvent
0.2(x+y) =x
y=0.8x
y=w/mol wt of benzene
w=0.8x(78)
Molality=x/0.8x(72)
X=0.0160